I am trying to solve by addition but my answers are wrong. Can you help me solve this equation by addition? I tried multiplying the first equation by 2 but it didn't work.
2x-4y=11
-4x+8y=-22

Answer by Alan3354(69443)

I am trying to solve by addition but my answers are wrong. Can you help me solve this equation by addition? I tried multiplying the first equation by 2 but it didn't work.
2x-4y=11
-4x+8y=-22
=====================
The 2nd eqn is -2 times the 1st.
--> the same eqn twice
Dependent, no unique solution.